# Debt, Growth, and the Carbon Lock-In

### v1

@@ Acknowledgments
We are grateful to Eric Galbraith, Julia Steinberger, Nico Wunderling, Jeff Althouse, and André de Palma for their insightful feedback and discussions, which greatly strengthened this work. S.M. acknowledges support from the 2025 FIR–PRI (Best Research Grant) for their PhD project.
Appendix A Gambling model and its connection to macroeconomic models for public debt
In this section, we present the gambling model studied in this paper, which is an extension of Kelly’s gambling model, and then, we explain how this extended Kelly’s model relates to known macroeconomic models for public debt.
A.1 Definition of the gambling model
We assume that at each time step $\tau$ (we use discrete time), a gambler can borrow an amount $B_{\tau}$ with interest rate $\rho_{\tau}$. The interest rate $\rho_{\tau}$ is the factor by which the amount borrowed is multiplied at each time step, in other words, the debt due to the amount borrowed at time $\tau^{\prime}$, is multiplied by $\rho_{\tau^{\prime}}^{\tau-\tau^{\prime}}$ if it has not been paid back at time $\tau$. In general, the interest rate is larger than 11 ($\rho>1$) so that the debt keeps increasing, and the borrower ends up paying more than the amount that was borrowed. Then the gross capital $C_{\tau+1}$, and debt $D_{\tau+1}$ after one step are (until bankruptcy occurs if $C_{\tau+1}\leq 0$):
$C_{0}$ $=B_{0}+C_{\text{ini}}=LC_{\text{ini}}$ (11)
$C_{\tau+1}$ $=\gamma_{\tau}C_{\tau}+B_{\tau+1}-\left(\sum_{\tau^{\prime}=0}^{\tau}\delta_{\tau,t_{p}(\tau^{\prime})}\right)D_{\tau+1}$ (12)
$D_{0}$ $=B_{0}=(L-1)C_{\text{ini}}$ (13)
where $\gamma_{\tau}$ is stochastic and depends on time (in a classical Kelly’s model, where $x$ represents the winner of time $\tau$, happening with a probability given by the probability vector $\boldsymbol{p}$, $\boldsymbol{b}$ is the gambling strategy and $\boldsymbol{r}$ is the inverse-odds vector, we would get that $\gamma_{\tau}=b_{x,\tau}/r_{x,\tau}$ which is a random variable which distribution depends on $\boldsymbol{p}$).. The first term in the first equation represents the standard evolution in Kelly’s model, namely the contribution of gambling at time $\tau$: if $\gamma_{\tau}>1$, the gross capital increases provided that there is no borrowing or payback at this time. The second term represents the amount borrowed at time $\tau$; it can be 0 if the borrower does not borrow at this step, or a positive amount. The third term represents the possible payback of the debt (debt clearing) at time $\tau+1$, and will either be equal to 0 (no payback), or to $D_{\tau+1}$. $C_{\text{ini}}$ is the amount invested initially by the gambler, and $B_{0}$ the initial amount borrowed. The leverage is defined as $L=(B_{0}+C_{\text{ini}})/C-i$ and quantifies the dependency of the gambler on an initial borrowing from the bank.
In addition $t_{p}(\tau^{\prime})$ indicates the time when the amount borrowed at $\tau^{\prime}$ is payed back, and the Heaviside function $\Theta$ function allows to know if the payback has occurred at time $\tau$ (indeed $\Theta(t_{p}(\tau^{\prime})-\tau)=0$ if $\tau>t_{p}$, meaning that the debt is erased after it has been payed back). If a payback happens at time $\tau$, it means that $\exists\tau^{\prime},\tau=t_{p}(\tau^{\prime})$, and thus $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=1$, otherwise $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=0$.
As explained above, without debt, the gross capital cannot hit 0 in Kelly’s model, and bankruptcy never occurs. In our updated view of this model, the capital can become negative if the debt exceeds the capital when the debt is repaid. Therefore solvency means that the debt is always paid back at a time where the gross capital is larger ($C_{\tau}>D_{\tau+}$). We call $C_{\tau}-D_{\tau}$ the net capital, whose sign indicates if bankruptcy will happen in case of payback.
A.2 Single loan model (strategy A)
For strategy A, we assume that the gambler initially borrows money ($B_{0}>0$) but never borrows afterwards, $\forall\tau\geq 1,B_{\tau}=0$. This strategy is more like that of an individual taking out a loan to buy a house. In practice, the gambler can borrow $B_{0}$ and invest $C_{\text{ini}}$ from its personal capital, and a key quantity will be the leverage $L=(B_{0}+C_{\text{ini}})/C_{\text{ini}}$. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model with an initial investment of $B_{0}+C_{\text{ini}}$, and the debt evolves with an interest rate $\rho$
@@ A.2 Single loan model (strategy A)
Obviously, this amount can be either negative or positive, which could not happen within the purely multiplicative Kelly’s model. Bankruptcy occurs if $C_{t_{p}}$ is negative or equivalently
@@
$w(t_{p})=\frac{1}{t_{p}}\ln\left(\frac{\left<C_{t_{p}}\right>}{LC_{\text{ini}}}\right)=\frac{1}{t_{p}}\ln\left(E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}}\right),$ (19)
which is the average growth rate of capital relative to the initial capital $LC_{\text{ini}}$ just after the debt is repaid, assuming no bankruptcy. This way we can write that $C_{t_{p}}=LC_{\text{ini}}\exp(w(t_{p})t_{p})$. The larger this quantity is, the faster the capital will grow after payback. If the gambler is thinking in short terms, this is a quantity to maximize. Interestingly we can bound this quantity (because $(1-1/L)\rho^{t_{p}}>0$) :
$\forall t_{p},\ln(E)\geq w(t_{p})$ (20)
@@
A.3 Periodic borrowing model (strategy B)
In this section, we assume that the gambler borrows money at each time step, with a leverage $L_{\tau}$ (depending on time a priori), which we refer to as strategy B. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model, with a modified stochastic growth rate, and the debt still evolves with an interest rate $\rho$. The gambler might want to partially pay back the debt at each step, but we ignore this possibility for now. Before the debt is paid, we have the equations modified in the form
@@ A.3 Periodic borrowing model (strategy B)
This sets the average short term growth rate in this case
$w(t_{p})=\frac{1}{t_{p}}\ln\left(\rho^{t_{p}}\left(\left(\frac{E+(L-1)}{\rho}\right)^{t_{p}}-1-\frac{L-1}{\rho}\frac{1-\left(\frac{E+(L-1)}{\rho}\right)^{t_{p}}}{1-\frac{E+(L-1)}{\rho}}+\frac{1}{L}\right)\right)$ (32)
@@
A.4 Relation to public debt models
In this section, we explicitly relate the model for the public debt of [33] to the gambling model outlined above. For the public debt, one usually studies debt in terms of the debt-to-GDP ratio. We call $C_{\tau}$ the GDP at time $\tau$, which can be split in different sectors : consumption $U_{\tau}$, investments $I(\tau)$, government spending $G_{\tau}$, exports $X_{\tau}$ and imports $M_{\tau}$ (setting $X_{\tau}-M_{\tau}$ to be the net exports). The primary deficit represents the value added to the public debt (apart from interests), and is split as government spending minus taxation $T_{\tau}$.
@@ A.4 Relation to public debt models
A.5 Relation to the Keen model
The Keen model [16] is a model that describes the dynamics of macroeconomic variables. In that model, the equations for $C_{\tau}$ and $D_{\tau}$, take the following form in terms of the total growth rate $g_{\tau}$, the total interest rate $r_{\tau}$, the investment to output ratio $\kappa_{\tau}$, the wage to labor ratio $\omega_{\tau}$, and the output to labor ratio $a_{\tau}$:
$\begin{split}C_{\tau+1}-C_{\tau}&=g_{\tau}C_{\tau}\\ D_{\tau+1}-D_{\tau}&=r_{\tau}D_{\tau}+\left(\kappa_{\tau}+\frac{\omega_{\tau}}{a_{\tau}}-1\right)C_{\tau}.\end{split}$ (39)
@@ A.5 Relation to the Keen model
$\begin{split}\gamma_{\tau}&\sim g_{\tau}-\left(\kappa_{\tau}+\frac{\omega_{\tau}}{a_{\tau}}\right)\\ \rho_{\tau}&\sim 1+r_{\tau}\\ L_{\tau}&\sim\left(\kappa_{\tau}+\frac{w_{\tau}}{a_{\tau}}\right).\end{split}$ (40)
In practice $\omega_{\tau}/a_{\tau}$ is the wage to output ratio, and $\kappa_{\tau}$ represents the share of output dedicated to investments. Following Fig.5, this suggests that the yearly variation of $\kappa_{\tau}+\omega_{\tau}/a_{\tau}$ is weak. The stochasticity of our model is thus encapsulated in $g_{\tau}$. This also gives insights on the variables, $L_{\tau}$ accounts for investments and labour.
@@ B.2 Energetic cost of debt
On Table.4 we report the different values extracted from data, for different countries, averaged over the period 199820221998-2022. To obtain $L,\rho,W$, we use the following equations and data for $GDP$, debt, and deficit
@@
where $d_{\tau+1}$ is the deficit of the state during year $\tau$ for year $\tau+1$, and the fact that $W=\left<\gamma_{\tau}\right$>.
China US Denmark France
$\left<L_{\tau}\right$> $1.025\pm 0.023$ $1.051\pm 0.040$ $0.997\pm 0.029$ $1.040\pm 0.018$
$\left<\rho_{\tau}\right$> $1.125\pm 0.023$ $1.010\pm 0.051$ $1.014\pm 0.155$ $1.002\pm 0.091$
$W$ $088\pm 0.086 0.0$$18\pm 0.059 0.0$$32\pm 0.089 0.0$$020\pm 0.090$
Table 4: Values of the different parameters.
@@
where $\epsilon(C,\tau)$ are the emissions of investing $C$ at time $\tau$, and $r(C,\tau)$ is the total use in resource when investing $C$ at time $\tau$. For the US, we obtain Fig.7 with the coefficients of determination $R^{2}=0.82$ for GDP and $R^{2}=0.79$ for debt. To ensure the model is working, we also use it to model GDP and debt for other countries and estimate their associated emissions. For France, we get the coefficients of determination $R^{2}=0.53$ for GDP and $R^{2}=0.86$ for debt. For China, we get the coefficients of determination $R^{2}=0.91$ for GDP and $R^{2}=0.97$ for debt. For Denmark, we obtain coefficients of determination $R^{2}=0.62$ for GDP and a negative $R^{2}$ for debt, indicating that the model fails to capture the strong variations and changes in sign of leverage. We observe a greater discrepancy between predictions and data for debt in Denmark. This is because Denmark usually runs a surplus and partially repays the debt. Therefore, the predicted debt exceeds the actual debt, which usually decreases. We report the coefficients of determination and rescaled mean absolute errors on table.5.
@@
Table 5: Rescaled Mean Absolute Error MAEMAE and coefficient of determination $R^{2}$ between data and theory for GDP and debt, for different countries. $MAE_{X}/\left<X\right>=\left(\sum_{n=1}^{N_{m}}|X_{n}-X_{n}^{\text{simu}}|\right)/\left(\sum_{n=1}^{N_{m}}X_{n}\right)$ with $N_{m}$ the number of measurements and $R^{2}_{X}=\sum\sum_{n=1}^{N_{m}}\left((X_{n}-X_{n}^{\text{simu}})/(X_{n}-\left<X\right>)\right)^{2}$.
Another important point to study is the difference between consumption-based emissions, meaning that some countries may display a comparatively low carbon intensity by only accounting for territorial emissions. However, when accounting for exported emissions, the carbon intensity increases. This is illustrated with predictions using our model in Fig.6.
Figure 6: Comparing territorial emissions and consumption-based emissions using our model for predictions. Territorial emissions represent carbon emissions produced on the territory of a given country, while consumption-based counts emissions due to activities directly used by a given country. For France and Denmark, we observe that territorial emissions are lower than consumption-based. Predictions are performed by running the model with the distributions of values for leverage $L$, $W$, $\rho$, and carbon intensity over the period 1998-2022. Carbon intensities are assumed to be linear in time, with a constant decarbonation rate corresponding to that of the period 1998-2022 (which is different for territorial emissions and consumption-based emissions).
@@
We also realize a plot similar to that of Fig.3 in the main text, showing predictions of the model for various intrinsic growth rates on Fig.8.
Figure 8: Predictions for cumulative emissions depending on the long-term growth rate $W$ in different countries. We also used our stochastic model and assume a linearly decreasing carbon intensity (with different parameters for each country). We assume that the leverage remains the same in each country.
@@ B.3 Comparing path dependent intensities to yearly intensities
from which we can easily see that if $\tau\to I(\tau)$ is a monotonically decreasing function, $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ will always be larger than $I(\tau)$. $I(\tau)$ is a decreasing function in many countries due to changes in energy use, often referred to as innovation. We can then compute the evolution of the quantity , defining $S_{\tau}=\sum_{\tau^{\prime}=0}^{\tau}C_{\tau^{\prime}}$
$\begin{split}\mathfrak{I}_{\tau+1}\left(\{C_{\tau}\}\right)-\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)&=\frac{1}{S_{\tau}S_{\tau+1}}\left(\sum_{\tau^{\prime}=0}^{\tau+1}I(\tau^{\prime})S_{\tau}C_{\tau^{\prime}}-\sum_{\tau^{\prime}=0}^{\tau}I(\tau^{\prime})S_{\tau+1}C_{\tau^{\prime}}\right)\\ &=\frac{1}{S_{\tau}S_{\tau+1}}\left(I(\tau+1)C_{\tau+1}S_{\tau}+\sum_{\tau^{\prime}=0}^{\tau}I(\tau^{\prime})C_{\tau^{\prime}}\left(S_{\tau}-S_{\tau+1}\right)\right)\\ &=\frac{C_{\tau+1}}{S_{\tau+1}}\left(I(\tau+1)-\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)\right).\end{split}$ (46)
The last equation leads to the conclusion that the path dependent intensity is decreasing with time if and only if $I(\tau+1)<\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ that is if and only if
$\sum_{\tau^{\prime}=0}^{\tau}\left(I_{\tau^{\prime}}-I(\tau+1)\right)C_{\tau^{\prime}}>0.$ (47)
This means that as long as the yearly carbon intensity is a decreasing function, so is the path dependent intensity. For a non monotonic intensity, the years with highest $C_{\tau}$ will prevail in the sum. From this we can also deduce the variations of $\Delta_{\tau}:=\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)$,
@@
which also means that for $I(\tau)$ strictly decreasing, $\Delta_{\tau}$ is initially increasing driven by the first term $S_{\tau}\left(I(\tau)-I(\tau+1)\right)/S_{\tau+1}$. Then it starts decreasing for $S_{\tau}\left(I(\tau)-I(\tau+1)\right)<C_{\tau+1}\Delta_{\tau}$. In the deterministic limit, we approximate $C_{\tau}\sim e^{r\tau}$ where $r$ is the growth rate, and obtain that this condition is
$\Delta_{\tau}>\left(I(\tau)-I(\tau+1)\right)\frac{1-e^{-r(\tau+1)}}{e^{r}-1}.$ (49)
@@
$\Delta_{\tau+1}\sim\left(I(\tau)-I(\tau+1)+\Delta_{\tau}\right)e^{-r}.$ (50)
@@
$\Delta_{\tau}\sim-\frac{1}{e^{r}-1}\frac{dI}{d\tau},$ (51)
meaning that $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)\sim I(\tau)-(dI/d\tau)/(e^{r}-1)$. In particular if we assume a power law dependency for the carbon intensity of the form $I(\tau)=I_{0}-\eta\tau^{n}$, the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ goes as
$\Delta_{\tau}\sim\frac{\eta n\tau^{n-1}}{e^{r}-1},$ (52)
therefore increasing to $\infty$ for $n>1$ and decreasing to 0 for $n<1$. In practice it means that for fast decarbonization, early intense years matter more in the path dependent intensity.
@@
In addition, if we forget about stochasticity and assume exponential growth $C_{\tau}=LC_{\text{ini}}e^{r\tau}$, we get
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)=I_{0}-\frac{\eta e^{r}}{e^{r}-1}\left(\frac{e^{r\tau}(\tau+1)(e^{r}-1)}{e^{r(\tau+1)}-1}-1\right),$ (54)
@@
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)=\eta\left(\tau+\frac{e^{r}((\tau+1)e^{r\tau}-\tau e^{r(\tau+1)}-1))}{(e^{r}-1)(e^{r(\tau+1)}-1)}\right),$ (55)
which is positive and increasing towards $1/(e^{r}-1)$ provided that $r>0$ as can be seen on Fig.9. This means that path dependent intensities are always larger than yearly intensities whenever GDP is growing over time.
Figure 9: A Difference between the path dependent intensity $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and the yearly intensity $I(\tau)$ with a deterministic evolution of $C_{\tau}$. The path-dependent intensity is systematically larger than the yearly intensity for $r>0$ and increases with time. It eventually reaches a strictly positive limit. B Effect of stochasticity on the difference between the path-dependent intensity $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and the yearly intensity $I(\tau)$. Increasing fluctuations modify the difference between the two quantities. C $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)$ follows a power law $n\eta\tau^{n-1}/(e^{r}-1)$ after a first transient phase, depending on the decarbonization rate.

### v2

@@ Acknowledgments
We are grateful to Eric Galbraith, Julia Steinberger, Jeff Althouse, and André de Palma for their insightful feedback and discussions, which greatly strengthened this work. S.M. acknowledges support from the 2025 FIR–PRI (Best Research Grant) for her PhD project.
Appendix A A stochastic macroeconomic credit-risk model
In this section, we provide more details on the stochastic credit-risk model studied in the section on Methods in the main text. There are two main variants of the model depending whether there is a single load (strategy A) or periodic borrowing (model B). Then, we also explain how this model relates to known macroeconomic models. In the next section, we provide more details on the data analysis and on the evaluation of carbon emissions.
A.1 Single loan model (strategy A)
For strategy A, we assume that the gambler initially borrows money ($B_{0}>0$) but never borrows afterwards, $\forall\tau\geq 1,B_{\tau}=0$. This strategy is more like that of an individual taking out a loan to buy a house. In practice, the gambler can borrow $B_{0}$ and invest $C_{\text{ini}}$ from its personal capital, and a key quantity will be the leverage $L=(B_{0}+C_{\text{ini}})/C_{\text{ini}}$. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model with an initial investment of $B_{0}+C_{\text{ini}}$, and the debt evolves with an interest rate $\rho-1$
@@ A.1 Single loan model (strategy A)
Bankruptcy occurs if $C_{t_{p}}$ is negative or equivalently
@@
$\omega(t_{p})=\frac{1}{t_{p}}\ln\left(\frac{\left<C_{t_{p}}\right>}{LC_{\text{ini}}}\right)=\frac{1}{t_{p}}\ln\left(E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}}\right),$ (19)
which is the average growth rate of capital relative to the initial capital $LC_{\text{ini}}$ just after the debt is repaid, assuming no bankruptcy. This way we can write that $C_{t_{p}}=LC_{\text{ini}}\exp(\omega(t_{p})t_{p})$. The larger this quantity is, the faster the capital will grow after payback. If the gambler is thinking in short-terms, this is a quantity to maximize. Interestingly we can bound this quantity (because $(1-1/L)\rho^{t_{p}}>0$) :
$\forall t_{p},\ln(E)\geq\omega(t_{p})$ (20)
@@
A.2 Periodic borrowing model (strategy B)
In this section, we assume that the gambler borrows money at each time step, with a leverage $L_{\tau}$ (depending on time a priori), which we refer to as strategy B. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model, with a modified stochastic growth rate, and the debt still evolves with an interest rate $\rho-1$. The gambler might want to partially pay back the debt at each step, but we ignore this possibility for now. Before the debt is paid, we have the equations modified in the form
@@ A.2 Periodic borrowing model (strategy B)
This sets the average short-term growth rate in this case
$\omega(t_{p})=\frac{1}{t_{p}}\ln\left(\rho^{t_{p}}\left(\left(\frac{E+(L-1)}{\rho}\right)^{t_{p}}-1-\frac{L-1}{\rho}\frac{1-\left(\frac{E+(L-1)}{\rho}\right)^{t_{p}}}{1-\frac{E+(L-1)}{\rho}}+\frac{1}{L}\right)\right)$ (32)
@@
A.3 Relation to public debt models
In this section, we explicitly relate the model for the public debt of [21] to the gambling model outlined above. For the public debt, one usually studies debt in terms of the debt-to-GDP ratio. We call $C_{\tau}$ the GDP at time $\tau$, which can be split in different sectors [22, 21]: consumption $U_{\tau}$, investments $I(\tau)$, government spending $G_{\tau}$, exports $X_{\tau}$ and imports $M_{\tau}$ (setting $X_{\tau}-M_{\tau}$ to be the net exports). The primary deficit represents the value added to the public debt (apart from interests), and is split as government spending minus taxation $T_{\tau}$.
@@ A.3 Relation to public debt models
A.4 Relation to the Keen model
The Keen model [3, 4, 6, 7] is a model inspired from Minsky, that describes the dynamics of macroeconomic variables. In that model, the equations for $C_{\tau}$ and $D_{\tau}$, take the following form in terms of the total growth rate $g_{\tau}$, the total interest rate $r_{\tau}$, the investment to output ratio $\kappa_{\tau}$, the wage to labor ratio $w_{\tau}$, and the output to labor ratio $a_{\tau}$:
$\begin{split}C_{\tau+1}-C_{\tau}&=g_{\tau}C_{\tau}\\ D_{\tau+1}-D_{\tau}&=r_{\tau}D_{\tau}+\left(\kappa_{\tau}+\frac{w_{\tau}}{a_{\tau}}-1\right)C_{\tau}.\end{split}$ (39)
@@ A.4 Relation to the Keen model
$\begin{split}\gamma_{\tau}&\sim g_{\tau}-\left(\kappa_{\tau}+\frac{w_{\tau}}{a_{\tau}}\right)\\ \rho_{\tau}&\sim 1+r_{\tau}\\ L_{\tau}&\sim\left(\kappa_{\tau}+\frac{w_{\tau}}{a_{\tau}}\right).\end{split}$ (40)
In practice $w_{\tau}/a_{\tau}$ is the wage to output ratio, and $\kappa_{\tau}$ represents the share of output dedicated to investments. Following Fig.5, this suggests that the yearly variation of $\kappa_{\tau}+w_{\tau}/a_{\tau}$ is weak. The stochasticity of our model is thus encapsulated in $g_{\tau}$ (which is typically deterministic in the Keen model [3, 7]). This also gives insights on the variables, $L_{\tau}$ accounts for investments and labour.
@@ B.2 Energetic cost of debt
On Table.1 of the main text we report the different values extracted from data, for different countries, averaged over the period 199820221998-2022. To obtain $L,\rho,W$, we use the following equations and data for $GDP$, debt, and deficit
@@
where $d_{\tau+1}$ is the deficit of the state during year $\tau$ for year $\tau+1$, and the fact that $W=\left<\ln(\gamma_{\tau})\right$> (we use the same definition as in model A).
@@
where $\epsilon(C,\tau)$ are the emissions of investing $C$ at time $\tau$, and $r(C,\tau)$ is the total use in resource when investing $C$ at time $\tau$. For the US, we obtain Fig.7 with the coefficients of determination $R^{2}=0.82$ for GDP and $R^{2}=0.79$ for debt. To ensure the model is working, we also use it to model GDP and debt for other countries and estimate their associated emissions. For France, we get the coefficients of determination $R^{2}=0.53$ for GDP and $R^{2}=0.86$ for debt. For China, we get the coefficients of determination $R^{2}=0.91$ for GDP and $R^{2}=0.97$ for debt. For Denmark, we obtain coefficients of determination $R^{2}=0.62$ for GDP and a negative $R^{2}$ for debt, indicating that the model fails to capture the strong variations and changes in sign of leverage. We observe a greater discrepancy between predictions and data for debt in Denmark. This is because Denmark usually runs a surplus and partially repays the debt. Therefore, the predicted debt exceeds the actual debt, which usually decreases. We report the coefficients of determination and rescaled mean absolute errors in Table 2. Note that in this table, we used the notations:
$\frac{MAE_{X}}{\left<X\right>}=\frac{\sum_{n=1}^{N_{m}}|X_{n}-X_{n}^{\text{simu}}|}{\sum_{n=1}^{N_{m}}X_{n}},$ (45)
with $N_{m}$ the number of measurements and
$R^{2}_{X}=1-\frac{\sum_{n=1}^{N_{m}}(X_{n}-X_{n}^{\text{simu}})^{2}}{\sum_{n=1}^{N_{m}}(X_{n}-\left<X\right>)^{2}}.$ (46)
@@
Table 2: Rescaled Mean Absolute Error MAEMAE and coefficient of determination $R^{2}$ between data and theory for GDP and debt, for different countries.
It is important to understand the difference between territorial and consumption-based emissions. Territorial emissions represent carbon emissions produced on the territory of a given country, while consumption-based counts emissions due to activities directly used by a given country. Thus, some countries may display a comparatively low carbon intensity by only accounting for territorial emissions, like France or Denmark. However, when accounting for exported emissions, their cumulative emissions increase. This is illustrated with predictions using our model in Fig.6.
In Fig.8, we show the predictions of the model for various intrinsic growth rates, which carries a similar message as Fig.3 of the main text.
Figure 6: Comparing territorial emissions and consumption-based emissions using our model for predictions. For France and Denmark, we observe that territorial emissions are lower than consumption-based. Predictions are performed by running the model with the distributions of values for leverage $L$, $W$, $\rho$, and carbon intensity over the period 1998-2022. Carbon intensities are assumed to be linear in time, with a constant decarbonation rate corresponding to that of the period 1998-2022 (which is different for territorial emissions and consumption-based emissions).
@@
Figure 8: Predictions for cumulative emissions depending on the long-term growth rate $W$ in different countries. We used our stochastic credit-risk model and assume a linearly decreasing carbon intensity (with different parameters for each country). We assume that the leverage remains the same in each country.
@@ B.3 Comparing path dependent intensities to yearly intensities
from which we can easily see that if $\tau\to I(\tau)$ is a monotonically decreasing function, $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ will always be larger than $I(\tau)$. $I(\tau)$ is a decreasing function in many countries due to changes in energy use, often referred to as innovation. We can then compute the evolution of the quantity,
$\begin{split}\mathfrak{I}_{\tau+1}\left(\{C_{\tau}\}\right)-\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)&=\frac{1}{S_{\tau}S_{\tau+1}}\left(\sum_{\tau^{\prime}=0}^{\tau+1}I(\tau^{\prime})S_{\tau}C_{\tau^{\prime}}-\sum_{\tau^{\prime}=0}^{\tau}I(\tau^{\prime})S_{\tau+1}C_{\tau^{\prime}}\right)\\ &=\frac{1}{S_{\tau}S_{\tau+1}}\left(I(\tau+1)C_{\tau+1}S_{\tau}+\sum_{\tau^{\prime}=0}^{\tau}I(\tau^{\prime})C_{\tau^{\prime}}\left(S_{\tau}-S_{\tau+1}\right)\right)\\ &=\frac{C_{\tau+1}}{S_{\tau+1}}\left(I(\tau+1)-\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)\right),\end{split}$ (48)
where we have introduced the notation $S_{\tau}=\sum_{\tau^{\prime}=0}^{\tau}C_{\tau^{\prime}}$.
The last equation leads to the conclusion that the path-dependent intensity is decreasing with time if and only if $I(\tau+1)<\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ that is, if and only if
$\sum_{\tau^{\prime}=0}^{\tau}\left(I(\tau^{\prime})-I(\tau+1)\right)C_{\tau^{\prime}}>0.$ (49)
This means that, as long as the yearly carbon intensity is a decreasing function, the path-dependent intensity is as well. For a non-monotonic intensity, the years with the highest $C_{\tau}$ will prevail in the sum. From this we can also deduce the variations of $\Delta_{\tau}:=\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)$,
@@
which also means that for $I(\tau)$ strictly decreasing, $\Delta_{\tau}$ is initially increasing driven by the first term $S_{\tau}\left(I(\tau)-I(\tau+1)\right)/S_{\tau+1}$. Then it starts decreasing for $S_{\tau}\left(I(\tau)-I(\tau+1)\right)<C_{\tau+1}\Delta_{\tau}$. In the deterministic limit, we approximate $C_{\tau}\sim e^{W\tau}$ at large $\tau$, where $W$ is the growth rate. Thus, we obtain that this condition is
$\Delta_{\tau}>\left(I(\tau)-I(\tau+1)\right)\frac{1-e^{-W(\tau+1)}}{e^{W}-1}.$ (51)
@@
$\Delta_{\tau+1}\sim\left(I(\tau)-I(\tau+1)+\Delta_{\tau}\right)e^{-W}.$ (52)
@@
$\Delta_{\tau}\sim-\frac{1}{e^{W}-1}\frac{dI}{d\tau},$ (53)
meaning that $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)\sim I(\tau)-(dI/d\tau)/(e^{W}-1)$. In particular if we assume a power law dependency for the carbon intensity of the form $I(\tau)=I_{0}-\eta\tau^{n}$, the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ goes as
$\Delta_{\tau}\sim\frac{\eta n\tau^{n-1}}{e^{W}-1},$ (54)
therefore increasing to $\infty$ for $n>1$ and decreasing to 0 for $n<1$ as $\tau\to\infty$. In practice, it means that for fast decarbonization, early intense years matter more in the path-dependent intensity.
@@
In addition, if we forget about stochasticity and assume exponential growth $C_{\tau}=LC_{\text{ini}}e^{W\tau}$, we get
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)=I_{0}-\frac{\eta e^{W}}{e^{W}-1}\left(\frac{e^{W\tau}(\tau+1)(e^{W}-1)}{e^{W(\tau+1)}-1}-1\right),$ (56)
@@
$\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)=\eta\left(\tau+\frac{e^{W}((\tau+1)e^{W\tau}-\tau e^{W(\tau+1)}-1))}{(e^{W}-1)(e^{W(\tau+1)}-1)}\right),$ (57)
which is positive and increasing towards $1/(e^{W}-1)$ provided that $W>0$ as can be seen on Fig.9. This means that path-dependent intensities are always larger than yearly intensities whenever GDP is growing over time.
Figure 9: A Difference between the path dependent intensity $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and the yearly intensity $I(\tau)$ with a deterministic evolution of $C_{\tau}$. The path-dependent intensity is systematically larger than the yearly intensity for $W>0$ and increases with time. It eventually reaches a strictly positive limit. B Effect of fluctuations in the economy on the difference between the path-dependent intensity $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and the yearly intensity $I(\tau)$. C $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)$ follows a power law $n\eta\tau^{n-1}/(e^{W}-1)$ after a first transient phase, depending on the decarbonization rate.